# UFT: Unifying Fine-Tuning of SFT and RLHF/DPO/UNA through a Generalized Implicit Reward Function

### v1

@@ 3.3 Data Distribution in Training UFT
Model gsm8k truthful-qa winograde arc hellaswag mmlu average
@@ 4 Related Work
To simplify RLHF, DPO has been proposed to map the optimal policy and reward model into a single step, transforming the initial unstable RL into a binary cross-entropy problem [RSM+23]. DPOP mathematically proves that during DPO, the reward for desired responses decreases and proposes a maximum term to prevent this [PKD+24]. IPO discovered that under nearly deterministic conditions between desired and undesired responses, the effectiveness of the KL divergence constraint imposed by $\beta$ diminishes, potentially leading to overfitting, and proposed a new loss term to address this issue [ARP+23]. sDPO suggests dividing the dataset into splits and using these splits sequentially to align the model, achieving better performance than using the entire dataset at once [KKS+24]. Iterative DPO posits that LLMs can serve as both response generators and evaluators, allowing for continuous improvement [YPC+24, XLSW24]. TDPO introduces the idea of providing rewards for each token generation [RHPF24, ZLM+24].
@@
There have also been efforts to merge SFT with alignment. ORPO proposed a new loss function to increase the ratio of desired responses over undesired responses, achieving both SFT and alignment [HLT24]. PAFT suggested conducting SFT and alignment in parallel and merging them afterward [PWB+24]. However, ORPO’s reliance on pairwise datasets and its deteriorating performance compared to other SFT and alignment methods pose challenges. In contrast, PAFT requires training separate adaptors for SFT and alignment and merging them through sparsity, which is inefficient. Inspired by UNA’s achievements, we aim to unify SFT with alignment based on UNA’s principles to avoid catastrophic forgetting.
@@ 5 Conclusion
Despite the extensive pretraining of LLMs using trillions of tokens and billions of parameters, they still fall short of being fully useful. SFT enhances LLMs with the capability to answer prompts effectively. However, even with pretraining and SFT, undesired responses can still occur. Alignment techniques, including RLHF, DPO, KTO and UNA, can improve the quality of alignment. Nevertheless, these sequential SFT+alignment methods can lead to catastrophic forgetting, where the model loses capabilities gained in previous stages.
To address this issue, a Unified Fine-Tuning (UFT) approach has been proposed. UFT unifies SFT and alignment by utilizing a generalized implicit reward function, following the UNA work. Trained solely on instruction-tuning data, UFT outperforms SFT, which is attributed to the minimization of divergence from the pretrained model through KL divergence. By mixing instruction-tuning data with alignment data, UFT surpasses all three sequential SFT+alignment methods, mitigating catastrophic forgetting. Ultimately, we establish a unified fine-tuning framework that runs in parallel to pretraining.
@@ Appendix A SFT on 20k examples with Different LR
LR gsm8k truthful-qa winograde arc hellaswag mmlu average
@@ Appendix B UFT on 20k examples with Different LR and $\beta$
LR/ $\beta$ gsm8k truthful-qa winograde arc hellaswag mmlu average

### v2

@@ 3 Experiments
Model gsm8k truthful winograde arc hellaswag mmlu average
@@
Table 9: Impact of different distributions of instruction-tuning and alignment data on MT-Bench and Alpaca-eval
3.1 UFT vs. SFT
To begin with, we compare UFT with SFT on the same instruction-tuning dataset. The UltraChat dataset is utilized by unfolding one conversation into multiple training examples to faciliate the training of UFT. From the UltraChat dataset, 20k samples are selected and utilized for training. The Mistral 7B-v0.1 [JSM+23] and Qwen 32B [QY+25] are utilized as the base model, with low rank adaptation (LoRA) of $r=16$ [HSW+21]. For SFT, the learning rate of $1e^{-4}$ is the best, and for UFT, the learning rate of $3e^{-5}$ and $\beta$ of 0.01 is the best.
The fine-tuned models are tested on different tasks including 12 tasks on two HuggingFace Open LLM Leaderboards[BFH+23, FHL+24]. The new Open LLM Leaderboard includes 6 tasks: bbh (Suzgun et al., 2022), gpqa (Rein et al., 2023), mmlu-pro (Wang et al., 2024), musr (Sprague et al., 2024), ifeval (Zhou et al., 2023), and math-hard (Hendrycks et al., 2021). For all these tasks, the average scores are reported. Conversely, the old Open LLM Leaderboard comprises 6 different tasks: gsm8k (Cobbe et al., 2021), truthful (Lin et al., 2022), winograde (Sakaguchi et al., 2019), arc (Allen AI), hellaswag (Zellers et al., 2019), and mmlu (Hendrycks et al., 2021). In this work, the average match rate in gsm8k, mc2 in truthful, accuracy in winograde, acc-norm in arc, acc-norm in hellaswag, and accuracy in mmlu will be reported. Additionally, MT-Bench [ZCS+23] and Alpaca-eval [LZD+23] will be used to evaluate the model’s ability to generate text responses, rather than selecting from predefined candidate answers. Specifically, the Length-Controlled Win Rate (LC WR) will be documented for Alpaca-eval. Additionally, the average length of the outputs within this evaluation will be provided for comprehensive analysis.
As demonstrated in Table 1 and Table 2, Mistral+UFT surpasses Mistral+SFT in 8 out of 12 tasks. When considering the average performances across these leaderboards, Mistral+UFT consistently outperforms Mistral+SFT. This indicates that UFT enhances the capabilities of an LLM by maximizing the reward and minimizing the difference from the pretrained model. Furthermore, in the MT-Bench evaluation, Mistral+UFT outperforms Mistral+SFT, whereas in the Alpaca-eval, Mistral+SFT overshadows Mistral+UFT. From a generation capability standpoint, both UFT and SFT exhibit similar performance. Similar conclusion can be found for Qwen 32B. Qwen+UFT outperforms Qwen+SFT in 10 out of 14 tasks as shown in Table 4 and Table 5 and both MT-Bench and Alpaca-eval in Table 6 . As a result, when evaluating the performance on downstream tasks, UFT demonstrates superiority over SFT.
3.2 UFT vs SFT+Alignment
Another advantage of UFT lies in combining SFT and alignment into one stage to avoid performance degradation. In this stage, we utilize the HelpSteer2 dataset of 20k examples for alignment [WDD+24]. For UFT, the 20k examples from UltraChat and 20k examples from HelpSteer2 are merged and utilized for training. For comparison, the best performing SFT model of learning rate $1e^{-4}$ in the previous experiment is utilized for further fine-tuning using DPO, KTO and UNA.
The same tasks are utilized for evaluation. In the two HuggingFace open LLM leaderboards, Mistral+UFT outperforms all of Mistral+SFT+DPO, Mistral+SFT+KTO, and Mistral+SFT+UNA in 9 out of 12 tasks. In terms of average scores, Mistral+UFT surpasses all three sequential methods. Several aspects need further discussion. Firstly, the performance degradation problem is evident as the performances of Mistral+SFT+DPO, Mistral+SFT+KTO, and Mistral+SFT+UNA are worse than Mistral+SFT, a phenomenon also known as alignment tax. Additionally, we observe that Mistral+UFT shows significant improvements on ifeval, which tests the model’s capability of instruction-following, and truthful, which assesses the model’s alignment capabilities. These results indicate that UFT greatly enhances instruction-following and alignment capabilities, demonstrating its effectiveness. In terms of generation capability, Mistral+UFT outperforms the other three sequential methods and it does not seem to suffer from performance degradation, as it performs better than both Mistral+SFT and Mistral+UFT on instruction-tuning data alone. Moreover, Mistral+UFT does not bias towards long generation like the other three sequential methods, which is another advantage of UFT. Simiar observations can be found for Qwen 32B. Qwen+UFT outperforms Qwen+SFT+Alignment in 9 out of 14 tasks, especially the average of the new and old HuggingFace open LLM leaderboards as shown in Table. 4 and Table. 5. For generation capability, Qwen+UFT outperforms Qwen+SFT+Alignment on both MT-Bench and Alpaca-eval, though the benefits diminish compared with smaller Mistal 7B model. This means that the performance degradation caused by SFT and alignment is diminished with larger model size, and this is consistent the conclusions of previous works on RLHF [BJN+22].
3.3 Data Distribution in Training UFT
Several intriguing points will be discussed. As shown in Table 7 and Table 8, among the 12 tasks, one task, namely musr, shows a minor statistical decrease in performance. In contrast, four tasks-gpqa, gsm8k, winograde, and arc—exhibit statistically small improvements. Notably, two tasks, ifeval and truthful, demonstrate statistically significant improvements. Specifically, ifeval improves from 23.22 to around 44, and truthful improves from 42.58 to around 53. Significant advancements are observed in MT-Bench and Alpaca-eval, indicating a substantial enhancement in the generation capability of the LLM as shown in Table 9. This aligns with our expectations, as instruction-tuning data primarily aim to enhance the model’s instruction-following capabilities, which positively impacts ifeval, MT-Bench, and Alpaca-eval. Meanwhile, alignment data contribute to improvements in tasks like truthful. Other tasks remain largely unaffected due to the absence of relevant data, suggesting that incorporating more pertinent data could enhance their performance. When increasing the proportion of instruction-tuningg data, we observe a decrease in truthful performance from 56 to 50, indicating that a larger proportion of alignment data benefits the bias and ethical performance of the LLM. On the other hand, simply adding more instruction-tuning data does not improve the performance of ifeval, MT-Bench, and Alpaca-eval. Therefore, further investigation into the optimal ratio between instruction-tuning data and alignment data is warranted.
@@ 4 Related Work
To simplify RLHF, DPO has been proposed to map the optimal policy and reward model into a single step, transforming the initial unstable RL into a binary cross-entropy problem [RSM+23]. Several downstream studies have been conducted on DPO to expand and generalize its capabilities, such as different loss functions [PKD+24, ARP+23], iterative approaches [YPC+24, XLSW24], token DPO [RHPF24, ZLM+24] and stepwise DPO [KKS+24].
@@
There have also been efforts to merge SFT with alignment. ORPO proposed a new loss function to increase the ratio of desired responses over undesired responses, achieving both SFT and alignment [HLT24]. PAFT suggested conducting SFT and alignment in parallel and merging them afterward [PWB+24]. However, ORPO’s reliance on pairwise datasets and its deteriorating performance compared to other SFT and alignment methods pose challenges. In contrast, PAFT requires training separate adaptors for SFT and alignment and merging them through sparsity, which is inefficient. Inspired by UNA’s achievements, we aim to unify SFT with alignment based on UNA’s principles to avoid performance degradation.
@@ 5 Conclusion
After the pretraining stage, the sequential SFT+Alignment methods can lead to performance degradation, where the model loses capabilities gained in previous stages. To address this issue, the UFT approach has been proposed. UFT unifies SFT and alignment by utilizing a generalized implicit reward function, following the UNA work. Trained solely on instruction-tuning data, UFT outperforms SFT, which is attributed to the minimization of divergence from the pretrained model through KL divergence. By mixing instruction-tuning data with alignment data, UFT surpasses all three sequential SFT+Alignment methods, mitigating performance degradation. Ultimately, we establish a unified fine-tuning framework that runs in parallel to pretraining.
6 Limitation
This paper has a couple of limitations that warrant attention in future research. First, the study is restricted to English, and the multilingual capabilities of the approach should be thoroughly evaluated. Additionally, the current research primarily utilizes datasets such as Ultrachat and Helpsteer, which are academic in nature. To enhance the applicability of the findings, further testing on industrial datasets is recommended.
@@ Appendix A SFT on 20k examples with Different LR
LR gsm8k truthful winograde arc hellaswag mmlu average
@@ Appendix B UFT on 20k examples with Different LR and $\beta$
LR/ $\beta$ gsm8k truthful winograde arc hellaswag mmlu average